OpenStudy (anonymous):

In an experiment, a physics student lies on a lightweight plank supported by two scales, 2.5m apart. If the left scale reads 290N and the right scale shows 122N, find the students mass. In that same situation, find the distance from the center of her head to her center of mass.

OpenStudy (anonymous):

Ok, consider the student is lying in the centre. So, you have F=mg... Hence, 290/9.8=29.6, and 122/9.8=12.4. Therefore a total mass of 42kg? Then you need to do centre of mass: \[x_{CM}=\frac{m_{1}x_{1}+m_{2}x_{2}}{m_{1}+m_{2}}\] and remember that x2 is 2.5-x1